OpenStudy (anonymous):

Solve. 4x over 14 = 3x + 11 over 16.

OpenStudy (anonymous):

4x/14 = (3x+11)/6 2x/7 = x/2 + 11/6 3x/14 = -11/6 x=-77/9

OpenStudy (anonymous):

my choices are 2, 4, 7 and 9.

OpenStudy (anonymous):

Oh, sorry. What is the value of x?

jhonyy9 (jhonyy9):

4x 3x+11 ---- = ------- 14 16 2*2x 3x+11 ------- = ------- simplifie by 2 the left side 2*7 16 2x 3x+11 ----- = ------- now multiply cross 7 16 2x * 16 = (3x+11)*7 32x=21x+77 subtract from both sides 21x 32x-21x=21x-21x+77 11x=77 divide both sides by 11 x=77/11 x=7

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

Thank you so much! ♥

jhonyy9 (jhonyy9):

yw good luck bye